genetics help, are the multiple choice answers ok

Questions 49 and 50 refer to this description In a


Questions 49 and 50 refer to this description In a
Answer #1

Answers to questions 49 and 50.

Assuming the population to be in accordance with Hardy Weinberg equilibrium.

Relevant Equation: (p+q)2 = p2+2pq+q2 = 1.

Individuals (GG for homozygous grey.)Individuals (bb for homozygous black.)Individuals (Gb for heterozygous grey.)

p= frequency of allele G ,q = frequency of allele b.

p2= frequency of individual GG

q2= frequency of individual bb

2pq = frequency of individual Gb

Answer 49 is E.In the given population,q2= frequency of individual bb is 9%=0.09.Therefore, q is square root of 9% i.e, 0.3.

Answer 50 is C. As p+q=1.From above,p = 1 - q = 1- 0.3 = 0.7.Therefore p2= (0.7)2 =0.49= 49%.

Answers to questions 51 and 52

Answer 51 is A Incomplete dominance

In regard to anemia phenotype in normal oxygen conditions, owing to the blood cell shape, it exhibits incomplete dominance.

Answer 52 is B Codominance.

In regard to the hemoglobin itself, it exhibits Codominance.